OpenStudy (babyslapmafro):

I am asked to find the differential at dy for the following function. y=xcos2x Can someone please get me started on this problem, I do not understand what I am being asked. thanks...

OpenStudy (anonymous):

product rule for this one \[(fg)'=f'g+g'f\] with \[f(x)=x,f'(x)=1,g(x)=\cos(2x),g'(x)=-2\sin(2x)\]

OpenStudy (babyslapmafro):

I looked in the book and found this... dy=f'(x)dx But I do not know how this relates to my problem...

OpenStudy (anonymous):

product rule... d/dx (u.v) = u. d/dx (v) + v. d/dx (u)

OpenStudy (anonymous):

u= x and v = cos2x put in above equation..

OpenStudy (babyslapmafro):

Do I just do the product rule and then multiply the result of the product rule by dx

OpenStudy (anonymous):

you'll get.. x d/dx (cos2x) + cos2x d/dx (x)

OpenStudy (anonymous):

 can you solve it now..??